OpenStudy (anonymous):

Find the equation of the line that passes through (-1,2) and (5,-4).

OpenStudy (anonymous):

\[ \frac { y+4}{x-5}= \frac {2+4}{-1-5} \]

OpenStudy (anonymous):

\[ \frac { y+4}{x-5}= \frac {2+4}{-1-5}=-1 \]

OpenStudy (anonymous):

\[ y+4 = -(x-5)\\ y+4 = -x + 5\\ y= -x+1 \]